# When Measurement Mediates the Effect of Interest

### v2

@@ Abstract
Many health promotion strategies aim to improve reach into the target population and outcomes among those reached. For example, an HIV prevention strategy could expand the reach of risk screening and the delivery of biomedical prevention to persons with HIV risk. This setting creates a complex missing data problem: the strategy improves health outcomes directly and indirectly through expanded reach, while outcomes are only measured among those reached. To formally define the total causal effect in such settings, we use Counterfactual Strata Effects: causal estimands where the outcome is only relevant for a group whose membership is subject to missingness and/or impacted by the exposure. To identify and estimate the corresponding statistical estimand, we propose a novel extension of Two-Stage targeted minimum loss-based estimation (TMLE). Simulations demonstrate the practical performance of our approach as well as the limitations of existing approaches.
Key Words: cluster-randomized control trials; missing data; mediation; targeted minimum loss-based estimation; counterfactual strata effects:
@@ 1 Introduction
Missing data remain a persistent challenge in medical and public health research [1]. Persons with complete data on key variables often differ meaningfully from persons with missing data. Failure to account for these differences can lead to biased estimates and precision losses [2, 3, 4]. Common approaches to handle missing data include (1) a ”complete-case analysis”, which restricts to participants with the relevant variables measured, (2) adjustment for baseline factors (e.g., age and sex) through standardization (i.e., G-computation) or inverse weighting, and (3) multiple imputation [5, 6, 7]. These approaches are not directly applicable when the total effect is of interest and measurement mediates the causal effect.
Concretely, health screening is often the first step to initiating a new treatment or prevention product. For example, HIV testing is a prerequisite for initiation of antiretroviral therapy (ART) for persons with HIV and for initiation of pre-exposure prophylaxis (PrEP) for persons at risk of HIV acquisition. Similarly, blood pressure must be measured prior to starting antihypertensive medication, and LDL cholesterol must be measured prior to starting Statins. However, the coverage of health screening is never 100%, and the target population of our prevention/treatment efforts is generally unknown (e.g., [8, 9]). Therefore, implementation strategies to improve health outcomes often also aim to improve the reach of health screening. For example, a strategy could expand outreach and service delivery into the community through village health teams or peer-based approaches (e.g., [10, 11, 12]). Likewise, a strategy could improve screening and uptake by integrating services at the health clinic or by offering new services at a pharmacy (e.g., [13, 14]).
@@
In such settings, the strategy is often deployed at the group-level or induces changes at the group-level. Examples of groups include communities, pharmacies, clinics, and health systems. Also in such settings, health screening is a mediator; it is an intermediate variable on the causal pathway between the exposure (a health promotion strategy) and the outcome (treatment/prevention uptake) [15, 16, 17, 18, 19, 20]. Several mediated effects may be of potential interest, including controlled direct effects, natural direct and indirect effects, stochastic direct and indirect effects, and separable direct and indirect effects [17, 21, 22, 23, 24]. These estimands correspond to hypothetical interventions to change the distribution of the mediator and are reviewed in Rudolph et al. [25]. However, when studying a novel health promotion strategy, it is of primary interest to evaluate the total causal effect, prior to investigating mediating pathways.
@@
We present a general framework to evaluate total effects of strategies that simultaneously improve reach into the target population and outcomes among those reached. Specifically, we provide a framework to address the novel multilevel-mediation-missing data problem, arising because (1) the group-level intervention improves health outcomes directly and indirectly through expanded reach; (2) simultaneously, outcomes are only measured among those reached. First, we give an overview of the motivating study, a cluster randomized trial (CRT) to improve PrEP uptake in rural Kenya and Uganda. Next, we describe the data generation process with a hierarchical causal model and formally define the causal estimand with Counterfactual Strata Effects, occurring when the outcome is only relevant for a group whose membership is subject to missingness and/or impacted by the exposure [26, 27, 28, 29, 30]. For identification and estimation of the corresponding statistical estimand, we use Two-Stage targeted minimum loss-based estimation (TMLE), which separates control for missing data from effect evaluation [27]. Our simulation study evaluates the finite sample performance of our approach and compares with alternatives. We close with a discussion of extensions to more complex settings and areas for ongoing research. To match our motivating example, we focus on CRTs throughout.
@@ 2 Motivating Study
Alcohol use is a persistent risk factor for HIV acquisition [31, 32, 33, 34, 35, 36, 37, 38] . In sub-Saharan Africa, attending drinking venues (i.e., bars) as a patron or an employee has been associated with increased HIV risk [39, 40, 41]. OPAL (Outreach and Prevention at ALcohol venues) is an ongoing CRT whose primary objective is to evaluate the effectiveness of community-based, recruitment strategies on PrEP uptake among adults at drinking venues in rural Kenya and Uganda (NCT05862857). The units of randomization (i.e., clusters) are groups of nearby drinking venues. At these venues, patrons and employees receive recruitment cards offering multi-disease screening (intervention) or HIV-focused screening (control) at their local health clinic. When these cards are presented at the clinic, the participant is offered screening as well as PrEP initiation if interested and eligible. The primary research question is what would be the difference in PrEP uptake among persons with HIV risk if everyone received a multi-disease versus HIV-focused recruitment card.
Figure 1 provides a simplified schematic of OPAL and illustrates the multilevel-mediation-missing data problem inherent in studies aiming to expand reach into the target population and health outcomes among those reached. First, CRTs have a hierarchical data structure; individuals are nested in clusters. Second, risk screening mediates the effect on prevention use. There is a direct effect of the intervention on prevention use as well as an indirect effect through risk screening. Third, everyone has an “underlying” HIV risk status, which is only observed if they participate in screening. In other words, HIV risk status is missing unless they engage in screening, which is also a prerequisite for starting PrEP. Existing methods for missing data would adjust for differences between persons screened versus not screened. However, these methods would block the indirect effect and bias estimates of overall effectiveness towards the null. We now present our novel approach for estimating total effects, while addressing the multilevel-mediation-missing data problem, common to CRTs aiming to improve reach into the target population and health outcomes among persons reached.
Figure 1: Simplified causal graph for the OPAL trial to increase HIV risk screening and prevention use among persons with HIV risk: individuals are nested within clusters; HIV risk screening mediates the intervention effect, and HIV risk status is missing (unobserved) if an individual does not screen. New methods are needed to evaluate overall effectiveness, while accounting for the multilevel-mediation-missing data problem.
@@ 3 Methods
We first introduce the notation that will be used to formalize the data generating process, specify the causal estimand, and derive the statistical estimand. Throughout, clusters are indexed with $j=\{1,\ldots,J\}$, and individuals are indexed with $i=\{1,\ldots,N_{j}\}$. The number of individuals in a given cluster $N_{j}$ often varies. Cluster-level variables are indexed by super-script $c$ and shared by all individuals in that cluster. For the $j^{th}$ cluster, $E^{c}_{j}$ denotes the cluster-level baseline covariates, and $A^{c}_{j}=1$ denotes randomization to the intervention arm, while $A^{c}_{j}=0$ denotes randomization to the control arm. Bold font denotes the set of individual-level variables for a given cluster. For the $N_{j}$ individuals in the $j^{th}$ cluster, $\bm{W}_{j}$ is the matrix of individual-level characteristics at baseline:
@@
In OPAL, for example, $\bm{W}_{j}$ could include the age, sex, alcohol use, mobility, and socioeconomic status for the individuals in cluster $j$.
Asterisks denote underlying values of variables. In OPAL, $\bm{Y1}_{j}^{*}=(Y1_{ij}^{*}:i=1,...,N_{j})$ is the vector of underlying indicators of HIV risk for the $N_{j}$ individuals from cluster $j$. More generally, $Y1$ is an indicator of being in the target population of interest (e.g., having HIV for the outcome of ART initiation or having high blood pressure for the outcome of antihypertensive medication use). These indicators are only observed for individuals participating in health screening. Therefore, we define $\bm{\Delta}_{j}=(\Delta_{ij}:i=1,\ldots,N_{j}$) as the vector of measurement indicators for cluster $j$. In OPAL, the HIV risk status of given individual is observed if they are screened at the health center (i.e., if $\Delta$=1) and is missing otherwise (i.e., if $\Delta$=0). Then we define the observed indicators as $Y1_{ij}=\Delta_{ij}\times Y1_{ij}^{*}$ for individual $i$ in cluster $j$ and as $\bm{Y1}_{j}=(Y1_{ij}:i=1,\ldots,N_{j}$) for cluster $j$. In OPAL, $Y1=1$ if the individual screened at risk for HIV acquisition and is zero otherwise (i.e., screened but was not at risk, or did not screen). Finally, we define the outcome vector for cluster $j$ as $\bm{Y2}_{j}=(Y2_{ij}:i=1,\ldots,N_{j}$). In OPAL, $\bm{Y2}$ consists of indicators of PrEP initiation. To match the running example where PrEP initiation occurs at the health clinic, we assume complete measurement, but this can easily be relaxed. Altogether, we denote the observed data for individual $i$ in cluster $j$ as $O_{ij}=(E^{c}_{j},W_{ij},A^{c}_{j},\Delta_{ij},Y1_{ij},Y2_{ij})$.
@@ 3.1 Structural causal model
In this model, the values of the random variables on the left-hand side are determined by the structural equations $(f_{E^{c}},f_{\bm{W}},f_{A^{c}},f_{\bm{Y1}^{*}},f_{\bm{\Delta}},f_{\bm{Y1}},f_{\bm{Y2}})$, which are non-parametric, and specify each variable’s “parents”, including unmeasured factors $(U_{E^{c}},U_{\bm{W}},U_{A^{c}},U_{\bm{Y1}^{*}},U_{\bm{\Delta}},U_{\bm{Y2}})$. Due to randomization, the unmeasured factors contributing to the intervention assignment $U_{A^{c}}$ are independent of the others. The corresponding directed acyclic graph is given in Figure 2.
To reflect the OPAL study, we have made several exclusion restrictions. First, the trial arm $A^{c}$ is completely randomized; this assumption can easily be relaxed to reflect covariate-dependent randomization schemes. Second, there is no impact of the trial arm $A^{c}$ on underlying HIV risk status $\bm{Y1}^{*}$. In other words, we assume the intervention does not impact the target population of interest. We relax this and other assumptions in Section 5. Third, there is no effect of underlying risk status $\bm{Y1}^{*}$ on screening $\bm{\Delta}$; this assumption is needed for identification, as detailed below (Section 3.4). Finally, by definition, observed HIV risk status $\bm{Y1}$ is only a function of underlying risk status $\bm{Y1}^{*}$ and measurement $\bm{\Delta}$. Using the causal model, we can generate counterfactuals and formally define the causal estimand of interest, as described next.
Figure 2: A simplified hierarchical causal graph to illustrate the data generating process in a CRT where the intervention improves health screening and health outcomes (among those screened). Since the causal model is defined at the cluster-level, we drop the subscript $j$ throughout. For simplicity, we have omitted the unmeasured factors: $\bm{U}=(U_{E^{c}},U_{\bm{W}},U_{A^{c}},U_{\bm{Y1}^{*}},U_{\bm{\Delta}},U_{\bm{Y2}})$.
@@ 3.2 Counterfactual Strata Effects
Recall our goal is to estimate the total effect: a contrast in the expected counterfactual outcome under the intervention versus the expected counterfactual outcome under the control. In OPAL, for example, our goal is to evaluate the expected difference in the counterfactual probability of starting PrEP for persons with HIV risk if all clusters received multi-disease recruitment cards ($A^{c}=1$) versus HIV-only recruitment cards ($A^{c}=0$). Our causal effect of interest is formally defined as
@@
is the counterfactual probability of starting PrEP under trial arm $a^{c}$, among persons at risk of HIV acquisition. Of course, we could consider alternative contrasts (e.g., ratios) and alternative summaries (e.g., weighting by cluster size) [44]. Crucially, our causal parameter $\Psi^{*}$ is a summary of counterfactual outcomes under the two strategies of interest, but does not modify measurement. In other words, $\Psi^{*}$ captures the direct and indirect effects of the intervention in the target population without enforcing 100% compliance on screening. Therefore, $\Psi^{*}$ is an intent-to-treat parameter.
@@
Due to the conditioning on underlying variables, causal parameters of this type have been termed “Counterfactual Strata Effects” [30, 29]. Counterfactual Strata Effects are causal estimands where the outcome is only relevant for a group whose membership is subject to missingness and/or impacted by the exposure. In the running example, PrEP uptake is only relevant for persons who are at risk of HIV acquisition, and the population with HIV risk is generally unknown. Counterfactual Strata Effects have previously been applied to study intervention effects on population-level HIV viral suppression (the proportion of all persons with HIV who are suppressing viral replication) and the incidence of latent tuberculosis infection (the proportion of persons at risk of tuberculosis who acquire it) [45, 27, 46, 47, 26, 28, 48, 49]. In both examples, the outcome and the target population are subject to missingness, and the study intervention may have had an impact on the target population. (See extensions in Section 5.)
@@
Counterfactual Strata Effects are related but distinct from Principal Strata Effects [50, 51]. In the latter, effects are defined conditional on latent classes where a post-baseline variable is used to cross-classify participants by joint potential values of that variable under each exposure level. Examples include average causal effects among compliers, defiers, always takers, or never takers. In OPAL, for example, we could estimate the average treatment effect among the latent subgroup of persons who would be always be screened for PrEP eligibility regardless of the recruitment card that they received. However, our goal is to evaluate the overall effectiveness of the health strategy among the entire target population.
Counterfactual Strata Effects are also related but distinct from the framework proposed by Young and colleagues for time-to-event outcomes with competing events [23, 24, 52, 53]. Specifically, when the competing event is a censoring event, contrasts of the expected counterfactual outcomes define a controlled direct effect: the impact if all censoring had been prevented. To understand how the competing event mediates the causal effect in such settings, Young and colleagues have also proposed separable direct and indirect effects. However, our goal remains to evaluate the total effect of the health promotion strategy among persons in the target population, including those not reached for health screening. We now propose developing, evaluating, and applying Two-Stage TMLE for this effect.

### v3

# A causal framework for evaluating the total effect of strategies aiming to expand screening and to improve outcomes

@@ Abstract
For many health conditions, there are highly efficacious treatment and prevention products. Maximizing their impact requires strategies that improve the reach of health screening in order to establish who could benefit. For example, HIV prevention strategies aim to expand risk screening and to improve uptake of pre-exposure prophylaxis (PrEP) among those experiencing risk. Often, these strategies induce changes at the group-level (e.g., health clinics or communities) and are evaluated through cluster randomized trials. This scenario creates a complex, multilevel-mediation-missing data problem for the following reasons. First, the strategy is delivered at the cluster-level, while health screening and outcomes are at the individual-level. Second, the strategy improves health outcomes directly and indirectly through improved health screening. Third, everyone has an “underlying” status, which is only observed among those screened. To formally define the total effect in such settings, we use Counterfactual Strata Effects: causal estimands where the outcome is only relevant for a group whose membership is subject to missingness and/or impacted by the exposure of interest. To identify and estimate the corresponding statistical estimand, we propose a novel extension of Two-Stage targeted minimum loss-based estimation (TMLE). Simulations demonstrate the practical performance of our approach as well as the limitations of existing approaches.
footnotemark: footnotetext: Abbreviations: ART, Antiretroviral Therapy; CRTs, Cluster Randomized Trials; PrEP, Pre-Exposure Prophylaxis; TMLE, Targeted Minimum Loss-based Estimation.
Key Words: counterfactual strata effects; cluster randomized trials; group randomized trials; measurement; mediation; missing data; screening; targeted minimum loss-based estimation
@@ 1 Introduction
Health screening is the first step to initiating a new treatment or prevention product. For example, HIV testing is a prerequisite for initiation of antiretroviral therapy (ART) for persons with HIV and for initiation of pre-exposure prophylaxis (PrEP) for persons at risk of HIV acquisition. Similarly, blood pressure must be measured prior to starting antihypertensive medication, and LDL cholesterol must be measured prior to starting statins. Thus, status awareness is the first pillar in the Care Cascade: the series of steps from reaching persons in need through their consistent use of treatment/prevention [62, 61, 60, 63].
The term “cascade” highlights the drop-offs occurring at each step of the process. Indeed, the coverage of health screening is rarely 100%. As a result, the focus population of our treatment/prevention efforts is generally unknown (i.e., subject to missingness).
To improve health at a population-level, we need strategies also improve the reach of health screening. Such a strategy could expand outreach and service delivery into the community through village health teams or peer-based approaches (e.g., [55, 27, 24]). Alternatively, a strategy could improve screening and uptake by integrating services at the health clinic or by offering new services at a pharmacy (e.g., [40, 26]). These strategies are often deployed at the group-level or induces changes at the group-level. Examples of groups include communities, pharmacies, clinics, and health systems.
@@
In these scenarios, health screening is a mediator; it is an intermediate variable on the causal pathway between the exposure (the strategy) and the outcome (uptake of treatment/prevention) [19, 44, 13, 32, 46, 41]. Health screening can also be considered to be a competing event, because the outcome can only occur if the person participates in screening [2, 70]. Several mediated effects may be of potential interest. Examples include controlled direct effects, natural direct and indirect effects, stochastic direct and indirect effects, and separable direct and indirect effects [13, 52, 68, 70, 56]. These estimands correspond to hypothetical interventions to change the distribution of the mediator, and many are reviewed in Rudolph et al. [51]. However, when studying a novel health strategy, it is of primary interest to evaluate its total causal effect, which includes the direct and indirect effects, prior to investigating mediating pathways.
Additionally, in these settings, health screening — the mediator — is the key measurement indicator. All persons have an “underlying” health status. For example, someone is either at risk of HIV acquisition or not. However, their status is only observed if they participate in screening. Furthermore, the strategy aims to improve outcomes among everyone in the focus population of interest (e.g., all persons with HIV risk) — regardless of whether they participate in screening. In other words, we are interested in the strategy’s effect without an additional intervention to enforce screening (i.e., eliminate the competing event) [70]. Therefore, we cannot simply evaluate the strategy’s effect among persons who screened. Additionally, standard regression-based approaches to account for differences between persons screened versus missed are not directly applicable, because they block the indirect effect and bias estimates of the total effect towards the null.
A new approach is needed to estimate the total effect at a population-level, while rigorously accounting for differences between those measured and those missed.
@@
In the following, we provide a novel framework to address the multilevel-mediation-missing data problem, which commonly arises when public health strategies aim to improve both health screening and health outcomes. First, we give an overview of OPAL, our motivating study, which was designed to improve PrEP uptake among persons with HIV risk in Kenya and Uganda. Next, we describe the data generating process with a hierarchical causal model [42, 5]. We formally define the causal estimand with Counterfactual Strata Effects, occurring when the outcome is only relevant for a group whose membership is subject to missingness and/or impacted by the exposure [9, 6, 4, 38, 21, 45]. For identification and estimation of the corresponding statistical estimand, we use Two-Stage targeted minimum loss-based estimation (TMLE), which separates control for missing data from effect evaluation [4, 38]. Our simulation study evaluates the finite sample performance of our approach and compares with common alternatives, including Single-Stage approaches. We close with a discussion of extensions to more complex settings and areas for ongoing research.
@@ 2 Motivating Study
Alcohol use is a persistent risk factor for HIV acquisition [28, 39, 18]. In sub-Saharan Africa, attending alcohol-serving venues as a patron or an employee has been associated with increased HIV risk [35, 69, 12, 31, 10].
OPAL (Outreach and Prevention at ALcohol venues) is an ongoing cluster randomized trial (CRT) whose primary objective is to evaluate the effectiveness of community-based, recruitment strategies on PrEP uptake among adults at alcohol-serving venues in rural Kenya and Uganda (NCT05862857). The units of randomization (i.e., clusters) are groups of nearby venues. At these venues, patrons and employees receive recruitment cards offering multi-disease screening (intervention) or HIV-focused screening (control) at their local health clinic. When these cards are presented at the clinic, the participant is offered screening as well as PrEP initiation if interested and eligible. The primary research question is what would be the difference in PrEP uptake among persons with HIV risk if all venues received multi-disease versus HIV-focused recruitment cards.
Figure 1 provides a simplified schematic of OPAL and illustrates the multilevel-mediation-missing data problem inherent in studies aiming to expand health screening and improve health outcomes. First, CRTs have a hierarchical data structure; individuals are nested in clusters. Second, risk screening mediates the effect on PrEP initiation. There is a direct effect of the intervention strategy on PrEP uptake as well as an indirect effect through risk screening. Third, everyone has an “underlying” HIV risk status, which is only observed if they participate in screening. In other words, HIV risk status is missing unless they engage in screening, and our interest in improving PrEP uptake among all persons with HIV risk. Furthermore, persons who screen likely differ meaningfully from those who do not. We now present our novel approach for estimating total effects, while addressing the multilevel-mediation-missing data problem.
Figure 1: Simplified causal graph for the OPAL trial to increase HIV risk screening and PrEP use among persons with HIV risk: individuals are nested within clusters; HIV risk screening mediates the intervention effect, and HIV risk status is missing (unobserved) if an individual does not screen. New methods are needed to evaluate overall effectiveness, while accounting for the multilevel-mediation-missing data problem.
@@ 3 Methods
We first introduce the notation that will be used to formalize the data generating process, specify the causal estimand, and derive the statistical estimand. Throughout, clusters are indexed with $j=\{1,\ldots,J\}$, and individuals are indexed with $i=\{1,\ldots,N_{j}\}$. The number of individuals in a given cluster $N_{j}$ often varies. Cluster-level variables are indexed by super-script $c$ and are shared by all individuals in that cluster. For the $j^{th}$ cluster, $E^{c}_{j}$ denotes the cluster-level baseline covariates, and $A^{c}_{j}=1$ denotes randomization to the intervention arm, while $A^{c}_{j}=0$ denotes randomization to the control arm. Bold font denotes the set of individual-level variables for a given cluster. For the $N_{j}$ individuals in the $j^{th}$ cluster, $\bm{W}_{j}$ is the matrix of individual-level characteristics at baseline:
@@
In OPAL, for example, $\bm{W}_{j}$ could include the age, sex, alcohol use, mobility, and socioeconomic status for the $N_{j}$ individuals in cluster $j$.
Asterisks denote underlying values of variables. Specifically, let $Y1_{ij}^{*}$ be in indicator that the individual $i$ in cluster $j$ is in the focus population. In OPAL, for example, $Y1^{*}$ indicates being at risk of HIV acquisition. We emphasize that $Y1^{*}$ is not the outcome. Other examples of $Y1^{*}$ could be indicators of having HIV for the outcome of ART initiation or having high blood pressure for the outcome of antihypertensive medication use. Let $\bm{Y1}_{j}^{*}=(Y1_{ij}^{*}:i=1,\ldots,N_{j})$ be the corresponding vector of indicators for cluster $j$. These indicators are only observed for individuals who participate in health screening. Therefore, we define $\bm{\Delta}_{j}=(\Delta_{ij}:i=1,\ldots,N_{j})$ as the vector of measurement indicators for cluster $j$. In OPAL, the HIV risk status of a given individual is observed if they are screened at the health center (i.e., if $\Delta$=1) and is missing otherwise (i.e., if $\Delta$=0). Then we define the observed indicators as $Y1_{ij}=\Delta_{ij}\times Y1_{ij}^{*}$ for individual $i$ in cluster $j$ and as $\bm{Y1}_{j}=(Y1_{ij}:i=1,\ldots,N_{j}$) for cluster $j$. In OPAL, $Y1=1$ if the individual screened at risk for HIV acquisition and is zero otherwise (i.e., did not screen or screened but was not at risk).
Finally, we define the outcome vector for cluster $j$ as $\bm{Y2}_{j}=(Y2_{ij}:i=1,\ldots,N_{j}$). In OPAL, $\bm{Y2}$ consists of indicators of PrEP initiation. Throughout, having the outcome ($Y2=1$) is only possible if the individual participates in screening and they are in the population of interest ($Y1=1$). Therefore, $Y2$ is deterministically zero if $Y1=0$. To match the running example where PrEP initiation occurs at the health clinic, we assume complete measurement of the outcome, but this can easily be relaxed. Altogether, we denote the observed data for individual $i$ in cluster $j$ as $O_{ij}=(E^{c}_{j},W_{ij},A^{c}_{j},\Delta_{ij},Y1_{ij},Y2_{ij})$.
@@ 3.1 Structural causal model
In this model, the values of the random variables on the left-hand side are determined by the structural equations $(f_{E^{c}},f_{\bm{W}},f_{A^{c}},f_{\bm{Y1}^{*}},f_{\bm{\Delta}},f_{\bm{Y1}},f_{\bm{Y2}})$, which specify each variable’s “parents”, including unmeasured factors $(U_{E^{c}},U_{\bm{W}},U_{A^{c}},U_{\bm{Y1}^{*}},U_{\bm{\Delta}},U_{\bm{Y2}})$. Due to randomization, the unmeasured factors contributing to the intervention assignment $U_{A^{c}}$ are independent of the others. The corresponding directed acyclic graph is given in Figure 2.
We have made several exclusion restrictions. First, the trial arm $A^{c}$ is completely randomized; this assumption can easily be relaxed to reflect covariate-dependent randomization schemes. Second, there is no impact of the trial arm $A^{c}$ on underlying status $\bm{Y1}^{*}$. In other words, we assume the trial arm does not change who is in the focus population (e.g., who is at risk of HIV acquisition). We relax this and other assumptions in Section 5. Third, there is no effect of underlying status $\bm{Y1}^{*}$ on screening $\bm{\Delta}$; this assumption is needed for identification, as detailed below (Section 3.4). Fourth, by definition, observed status $\bm{Y1}$ is only a function of underlying status $\bm{Y1}^{*}$ and measurement $\bm{\Delta}$. Finally, the only effect of screening ($\bm{\Delta})$ on the outcome ($\bm{Y2}$) is through establishing that the individual is in the focus population ($\bm{Y1}=1$). Using the causal model, we can generate counterfactuals and formally define the causal estimand of interest, as described next.
Figure 2: A simplified hierarchical causal graph to illustrate the data generating process in a CRT where the intervention strategy expands health screening and improves health outcomes. Since the causal model is defined at the cluster-level, we drop the subscript $j$ throughout. For simplicity, we have omitted the unmeasured factors: $\bm{U}=(U_{E^{c}},U_{\bm{W}},U_{A^{c}},U_{\bm{Y1}^{*}},U_{\bm{\Delta}},U_{\bm{Y2}})$.
@@ 3.2 Counterfactual Strata Effects
Recall our goal is to estimate the total effect: a contrast in the expected counterfactual outcome under the intervention versus the expected counterfactual outcome under the control. In OPAL, for example, our goal is to evaluate the expected difference in the counterfactual probability of starting PrEP for all persons with HIV risk if all clusters received multi-disease recruitment cards ($A^{c}=1$) versus HIV-only recruitment cards ($A^{c}=0$). Our causal effect of interest is formally defined as
@@
is the counterfactual probability of starting PrEP under trial arm $a^{c}$, among persons at risk of HIV acquisition. Of course, we could consider alternative contrasts (e.g., ratios) and alternative summaries (e.g., weighting by cluster size) [11]. Crucially, our causal parameter $\Psi^{*}$ is a summary of counterfactual outcomes under the two strategies of interest, but does not involve a hypothetical intervention to enforce measurement of status. In other words, $\Psi^{*}$ captures the direct and indirect effects of the intervention in the focus population without enforcing 100% compliance on screening. Therefore, $\Psi^{*}$ can be classified as an intent-to-treat parameter.
@@
Due to the conditioning on underlying variables, causal parameters of this type have been termed “Counterfactual Strata Effects” [45, 21]. Counterfactual Strata Effects are causal estimands where the outcome is only relevant for a group whose membership is subject to missingness and/or impacted by the exposure. In the running example, PrEP uptake is only relevant for persons who are at risk of HIV acquisition, and the population with HIV risk is generally unknown. Counterfactual Strata Effects have previously been applied to study intervention effects on population-level HIV viral suppression (the proportion of all persons with HIV who are suppressing viral replication) and the incidence of latent tuberculosis infection (the proportion of persons at risk of tuberculosis who acquire it) [9, 4, 43, 22, 6, 38, 1, 33]. In both examples, the outcome and focus population are subject to missingness, and the study intervention may have had an impact on focus population. (See extensions in Section 5.)
@@
Counterfactual Strata Effects are related but distinct from Principal Strata Effects [20, 16]. In the latter, effects are defined conditional on latent classes where a post-baseline variable is used to cross-classify participants by joint potential values of that variable under each exposure level. Examples include average causal effects among compliers, defiers, always takers, or never takers. In OPAL, for example, we could estimate the average treatment effect among the latent subgroup of persons who would be always be screened for PrEP eligibility regardless of the recruitment card that they received. However, our goal is to evaluate the overall effectiveness of the health strategy among the entire focus population: persons with HIV risk. Counterfactual Strata Effects are also related but distinct from the framework proposed by Young and colleagues for time-to-event outcomes with competing events [70, 56, 34, 25]. Throughout, health screening can be considered a competing event. Concretely, one cannot initiate PrEP without first screening for eligibility. While Young et al. study competing events occurring over time among a well-defined cohort of participants, we center on evaluating outcomes among a focus population that is subject to missingness. In other words, our goal is to evaluate the total effect of the public health strategy among all persons in the group of interest, including those not reached for health screening. We now propose developing, evaluating, and applying Two-Stage TMLE for this effect.